OpenStudy (anonymous):

What is the 7th term of the geometric sequence where a1 = 625 and a2 = -125?

OpenStudy (amistre64):

a2/a1 = r then use the formula of a geo seq to determine the nth term an = a1* r^(n-1)

OpenStudy (anonymous):

what? :/ that doesnt make sense

OpenStudy (amistre64):

it has to make sense; there isnt a simpler way to explain it :/

OpenStudy (anonymous):

625 * ... idk what r is. is that like, s?

OpenStudy (anonymous):

can you type out the formula with numbers?

OpenStudy (amistre64):

you know the value of a2 you know the value of a1 a2/a1 = r .... so you know that value of r with what is given

OpenStudy (amistre64):

since your nth term is the 7th term, your n value is 7 in the second case

OpenStudy (amistre64):

like "s"? convention tends to use "r" to define the common ratio with

OpenStudy (anonymous):

625 * 5^6 ?

OpenStudy (amistre64):

that has a slight error, but it works out the same :) (-5)^6 instead of (5)^6 ... but since the power is even they work out the same in this case :)

OpenStudy (amistre64):

so yes

OpenStudy (anonymous):

97565625 ?

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

big numbers are allowed :)